Circular No. 8169

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gr IN MCG –04-55-14 Further to IAUC 8168, M. Ganeshalingam and W. Li report the LO-

TOSS discovery of an apparent supernova on unfiltered KAIT images taken on July 27.5 (mag ∼ 16.3) and 28.5 UT (mag ∼ 16.4). SN 2003gr is located at α = 23^{h}38^{m}12^{s}.67, δ = −20^{o}46′56′′.0 (equinox 2000.0), which is 5′′.9 west and 0′′.8 north of the nucleus of MCG –04-55-14. A KAIT image taken on 2002 Oct. 29.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002lh A. Gal-Yam and D. Maoz, Tel Aviv University; F. Prada, Isaac Newton

Group of Telescopes, IAC; and P. Guhathakurta, University of California, Santa Cruz, report the discovery of a supernova (V ∼ 21) on unfiltered images taken by Maoz on 2002 June 9 with the 2.5-m Nordic Optical Telescope (+ ALFOSC) at La Palma. SN 2002lh was located at α = 14^{h}44^{m}20^{s}.38, δ = +31^{o}15′44′′.4 (equinox 2000.0), in the field of the galaxy cluster Abell 1961 (z = 0.232). SN 2002lh was invisible on images obtained in April 2002 (limiting mag V ∼ 24.0). A. V. Filippenko and R. Chornock, University of California, Berkeley; and G. Smith, University of California, Santa Cruz, report that inspection of a CCD spectrum (range 550–930 nm), obtained by Smith on 2002 June 16 with the Keck-I telescope (+ LRIS spectrograph), reveals the object to be a type-Ia supernova, then near maximum brightness. The redshift (measured from narrow nebular lines from the host galaxy) is 0.236, consistent with that of the cluster Abell 1961.

SUPERNOVA 2003gj IN NGC 7017 R. J. Foley, L. Desroches, M. Ganeshalingam, W. Li, R. Chornock, and

A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 320–1000 nm) obtained on July 7 UT with the Shane 3-m telescope at Lick Observatory reveals that SN 2003gj (IAUC 8161) is probably of type Ia, ∼ 1 week past maximum brightness. However, its spectrum appears to be somewhat peculiar for this epoch, being dominated by strong emission lines near rest wavelengths of 550 nm and 645 nm, and a weaker one near 585 nm. Features blueward of 530 nm are very weak, as is the normally prominent O I 777.4-nm absorption. The spectrum is similar to that of the somewhat subluminous (unfiltered absolute magnitude at maximum ≈ −18 to −18.5) type-Ia supernovae 2003M and 2003ek at comparable epochs.

2003 July 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green